OpenStudy (anonymous):

\[\huge \int\limits \frac {e^{8/x}}{x^2} dx\] = ? + C (looking mainly for help with setup with tips & tricks; will give a medal :-) )

OpenStudy (anonymous):

Put \[u=\frac 8 x \]

OpenStudy (anonymous):

\[ du = -\frac 8 {x^2} dx \]

OpenStudy (anonymous):

\[ \int \frac{e^{8/x}}{x^2} \, dx= -\frac 1 8 \int e^u du=-\frac 1 8 e^u + C \]

OpenStudy (anonymous):

\[ \int \frac{e^{8/x}}{x^2} \, dx=-\frac{e^{8/x}}{8} + C \]

OpenStudy (anonymous):

Ah so a simple u substitution was thrown in there with the others... >_< Doh! Well ty for the help

OpenStudy (anonymous):

yw